Solve the exponential equation. Give the answer correct to 3 decimal places.
15,000 = 500e^03x

Answer by RAY100(1637)

15000=500e^3x
divide by 500
30=e^3x
ln both sides
ln30=3x
3.40=3x
x=1.1337
check 15,000=500e^(3*1.1337)=500(30)=15,000,,,ok